Can you help me with this problem?
7^3 divided by 7

Found 3 solutions by checkley77, Edwin McCravy, jhajha0830:
Answer by checkley77(12844)

7^3/7=7^(3-1)=7^2 ans.

Answer by Edwin McCravy(20054)
Answer by jhajha0830(1)